OpenStudy (anonymous):

what are the zeros for f(x)=2x^2-17?

OpenStudy (anonymous):

x = sqrt(17/2) and x = -sqrt(17/2)

OpenStudy (saifoo.khan):

\[-\sqrt{\frac{17}{2}}, \sqrt{\frac{17}{2}}\]